# On infinite versions of the prisoner problem

### v1

@@ Abstract
We investigate some versions of the famous 100 prisoner problem for the infinite case, where there are infinitely many prisoners and infinitely many boxes with labels. In this case, many questions can be asked about the admissible steps of the prisoners, the constraints they have to follow and also about the releasing conditions. We will present and analyze many cases. In the infinite case, the solutions and methods require mainly analysis rather than combinatorics.
@@ 2 Versions of the problem
We will denote the price of the $n^{th}$ box by $p_{n}$, and the amount that the $n^{th}$ prisoner gets by $a_{n}$, i.e. the strategy of the prisoners is the sequence $(a_{n})$. We require that $p_{n},a_{n}\geq 0$ and $\sum_{n=1}^{\infty}p_{n}=\sum_{n=1}^{\infty}a_{n}=1$.
@@ 2.1.1 Solutions
First let us analyze what ”there is no strategy” means. For a strategy $(a_{n})$, we require that for ALL finite-cycle-permutations of the labels, infinitely many prisoners find his label. A strategy depends on the price distribution only. Hence ”there is no strategy” is equivalent with that for every $(a_{n})$ there is a finite-cycle-permutation $\sigma:\mbox{$\mathbb{N}$}\to\mbox{$\mathbb{N}$}$ for which only finitely many prisoners will find his label. In other words, if the prisoners produce $(a_{n})$, then for that, the guard can make a finite-cycle-permutations of the labels such that only finitely many prisoners will find his label.
Let $(a_{n})$ be the strategy of the prisoners. Let us reorder $(a_{n})$ to become descending, i.e. find a permutation $\delta:\mbox{$\mathbb{N}$}\to\mbox{$\mathbb{N}$}$ such that $(a_{\delta(n)})$ is descending. To make to notation simpler, we can assume that $(a_{n})$ is descending, i.e. we solve the problem for descending $(a_{n})$ with the corresponding $(p_{n})$, and after solving it, apply $\delta^{-1}$ for both sequences to get back the original ones.
@@
    First he reorders $(a_{n})$ to become descending (as we did above).
@@
Clearly in each such cycle, the first member will not find his label by (3a). But as $(a_{n})$ is descending, none of the rest members in the cycle will. Hence after the first such cycle, no one will find his label. ∎
We can now merge theorems 2.3 and 2.6 into one.
@@
We can characterize those sequences for which this sum is infinite for all permutations.
@@
Recall from 2.5 that $\alpha_{n}=\delta(n)-n$, and $\delta(n)p_{n}=(n+\alpha_{n})p_{n}=np_{n}+\alpha_{n}p_{n}$. So in our case let
@@
Then all rest of the proof of 2.5 can be repeated thereafter. ∎
Analyzing the proofs of 2.5 and 2.8, we can realize that some parts can be applied to sequences with $\sum_{n=1}^{\infty}np_{\delta(n)}<\infty$ for a permutation $\delta:\mbox{$\mathbb{N}$}\to\mbox{$\mathbb{N}$}$.
@@
Apply the corresponding parts of theorems 2.5 and 2.8. ∎

### v2

@@ Abstract
We investigate some versions of the famous 100 prisoner problem for the infinite case, where there are infinitely many prisoners and infinitely many boxes with labels. In this case, many questions can be asked about the admissible steps of the prisoners, the constraints they have to follow and also about the releasing conditions. We will present and analyze many versions and cases. In the infinite case, the solutions and methods require mainly analysis rather than combinatorics.
@@ 2 Versions of the problem
Later, when we have a necessary and sufficient condition for the existence of a strategy, we will see that many of the above defining conditions can be weakened (see remarks 2.8, 2.10, 2.11, 2.12).
@@
We will denote the price of the $n^{th}$ box by $p_{n}$, and the amount that the $n^{th}$ prisoner gets by $a_{n}$, i.e. the strategy of the prisoners is the sequence $(a_{n})$. We require that $p_{n}\geq 0,\ a_{n}\geq 0$ and $\sum_{n=1}^{\infty}p_{n}=\sum_{n=1}^{\infty}a_{n}=1$.
@@ 2.1.1 Solutions
  Remark 2.3.
From the proof, we can see that theorem 2.2 remain valid if the prisoners get any (small) finite amount (not 1 as originally).
@@
First let us analyze what ”there is no strategy” means. For a strategy $(a_{n})$, we require that for ALL finite-cycle-permutations of the labels, infinitely many prisoners find his label. A strategy depends on the price distribution only. Hence ”there is no strategy” is equivalent with that for every $(a_{n})$ there is a finite-cycle-permutation $\sigma:\mbox{$\mathbb{N}$}\to\mbox{$\mathbb{N}$}$ for which only finitely many prisoners will find his label. In other words, if the prisoners produce $(a_{n})$, then for that, the guard can make a finite-cycle-permutation of the labels such that only finitely many prisoners will find his label.
Let $(a_{n})$ be the strategy of the prisoners.
If there are zeros among the elements of $(a_{n})$, then distribute amount 1 among them somehow (does not matter how). I.e. modify the values of such $a_{n}$-s by adding altogether 1 to them. E.g. if there are infinitely many zeros, then modify the zeros to $\frac{1}{2},\frac{1}{4},\frac{1}{8},\dots$. To simplify the notation, let us denote this new sequence by $(a_{n})$ as well. Clearly the sum changes to $\sum_{n=1}^{\infty}a_{n}=2$.
Let us reorder $(a_{n})$ to become descending, i.e. find a permutation ${\delta:\mbox{$\mathbb{N}$}\to\mbox{$\mathbb{N}$}}$ such that $(a_{\delta(n)})$ is descending. To make to notation simpler, we can assume that $(a_{n})$ is descending, i.e. we solve the problem for descending $(a_{n})$ with the corresponding $(p_{n})$, and after solving it, apply $\delta^{-1}$ for both sequences to get back the original ones.
@@
    First he changes the zeros in $(a_{n})$ by adding altogether 1 to them (as we did above).
@@
Clearly in each such cycle, the first member will not find his label by (4a) (if the original $a_{n}$ was 0, then it is trivial). But as $(a_{n})$ is descending, none of the rest members in the cycle will (again, if some member was originally 0, then it is trivial). Hence after the first such cycle, no one will find his label. ∎
  Remark 2.8.
From the proof, we can see that theorem 2.7 remains valid if the prisoners get any finite amount (not 1 as originally), because we just used it in inequality 1 and its consequence.
Together with remark 2.3, we can say that the answer (if there is a strategy) does not depend on how much (finite) amount the prisoners get altogether.∎
We can now merge theorems 2.4 and 2.7 into one.
@@
  Remark 2.10.
We can weaken the condition regarding how the guard is supposed to create his permutation. It can be readily seen that corollary 2.9 remains valid if we just require that the guard creates permutation such that it contains infinitely many finite cycles (and we allow that it also contains infinite cycles as well).∎
  Remark 2.11.
Moreover we can drop the ”finite-cycle-permutation” condition completely, if we also change how the box opening happens. When a prisoner enters the room, he shows his money to the guard and he points to a box, and he asks the guard to open all boxes belonging to the full cycle which contains that box, provided his money covers all boxes. If his money does not cover all, then he has to say which boxes he wants to open one-by-one. Again, in this case, corollary 2.9 remains valid.∎
  Remark 2.12.
As a consequence of remark 2.8, we can say that corollary 2.9 remains valid if we modify one more condition, namely, we allow the prisoners to send back the not used amount. Moreover everyone can send back all his amount. Because it acts like we just double the original full amount (which is 1).∎
We can characterize those sequences for which this peculiar sum in corollary 2.9 is infinite for all permutations. First, we characterize it for positive sequences.
@@
Recall from 2.6 that $\alpha_{n}=\delta(n)-n$, and $\delta(n)p_{n}=(n+\alpha_{n})p_{n}=np_{n}+\alpha_{n}p_{n}$. So in our case let
@@
Then all rest of the proof of 2.6 can be repeated thereafter. ∎
Before the general case, we need a little preparation.
  Definition 2.14.
Let $(a_{n})$ be a non-negative sequence. We call $(a_{n})$ quasi-descending, if $0<a_{n},\ n<m$ implies that $a_{m}\leq a_{n}$ holds.∎
  Proposition 2.15.
Let $(a_{n})$ be a non-negative sequence. Then $(a_{n})$ can be reordered to become quasi-descending.
Proof.
Let us order the strictly positive elements in a descending order, then add the remaining 0 elements somehow. ∎
  Theorem 2.16.
Let $(p_{n})$ be a sequence. Let $(q_{n})$ be the sequence that we get when we omit the zero terms from $(p_{n})$. Then the following statements are equivalent.
  1. 1.
    $\sum_{n=1}^{\infty}np_{\delta(n)}=\infty$ holds for every permutation $\delta:\mbox{$\mathbb{N}$}\to\mbox{$\mathbb{N}$}$.
  2. 2.
    $\sum_{n=1}^{\infty}np_{\delta(n)}=\infty$ holds when $\delta$ is a permutation which orders the sequence to be quasi-descending.
  3. 3.
    ${\sum_{n=1}^{\infty}nq_{\delta(n)}=\infty}$ holds for every permutation $\delta:\mbox{$\mathbb{N}$}\to\mbox{$\mathbb{N}$}$.
  4. 4.
    ${\sum_{n=1}^{\infty}nq_{\delta(n)}=\infty}$ holds when $\delta$ is a permutation which orders the sequence $(q_{n})$ to be descending.
Proof.
If $p_{n}>0$, then let $\alpha(n)$ denote the index of the number $p_{n}$ in the sequence $(q_{n})$, in other words let $q_{\alpha(n)}=p_{n}$ for $p_{n}>0$. Clearly $\alpha:\mbox{$\mathbb{N}$}\to\mbox{$\mathbb{N}$}$ is injective and descending.
The equivalence of 3 and 4 is in theorem 2.13. Clearly the implications $1\Rightarrow 2$ and $3\Rightarrow 4$ are obvious.
$3\Rightarrow 1$: If there is a permutation $\delta$ on $(p_{n})$, then it naturally determines a permutation $\sigma$ on the positive elements of $(p_{n})$ i.e. on $(q_{n})$. But the sum gets smaller: $\infty=\sum_{n=1}^{\infty}nq_{\sigma(n)}\leq\sum_{n=1}^{\infty}np_{\delta(n)}$.
$4\Rightarrow 2$: The same argument works with noting that a quasi-descending ordering on $(p_{n})$ is a descending ordering on $(q_{n})$.
$1\Rightarrow 3$: Let $\delta:\mbox{$\mathbb{N}$}\to\mbox{$\mathbb{N}$}$ be a permutation (applied for $(q_{n})$). We define a permutation $\sigma$ of $(p_{n})$. Let $\sigma:\mbox{$\mathbb{N}$}\to\mbox{$\mathbb{N}$}$ be the following. If $p_{n}>0$, then let $\sigma(n)=2\delta(\alpha(n))\ \mbox{$(n\in\mbox{$\mathbb{N}$})$}$, i.e. we do exactly the same like $\delta$ does, but on the set of even numbers only. For odd numbers, we somehow arrange the zeros of $(p_{n})$, that is how we define $\sigma$. Clearly we get that $\infty=\sum_{n=1}^{\infty}np_{\sigma(n)}=2\sum_{n=1}^{\infty}nq_{\delta(n)}$. Hence $\sum_{n=1}^{\infty}nq_{\delta(n)}=\infty$.
$2\Rightarrow 4$: The previous argument works with noting that $\sigma$ is a quasi-descending ordering of $(p_{n})$ if $\delta$ is a descending ordering of $(q_{n})$. ∎
Now we can simply derive if we omit the zeros in $(p_{n})$, then we get an equivalent sequence regarding strategy.
  Corollary 2.17.
Let $(p_{n})$ be given. Let us denote by $(q_{n})$ the sequence that we get when we omit the zeros from $(p_{n})$. Then there is strategy for $(p_{n})$ if and only if there is one for $(q_{n})$.
Proof.
Corollary 2.9 and theorem 2.16. ∎
Analyzing the proofs of 2.6 and 2.13, we can realize that some parts can be applied to sequences with $\sum_{n=1}^{\infty}np_{\delta(n)}<\infty$ for a permutation $\delta:\mbox{$\mathbb{N}$}\to\mbox{$\mathbb{N}$}$.
@@
Apply the corresponding parts of theorems 2.6 and 2.13. ∎
@@
2.4 Version 1.d
It is a modification of Version 1.b (see section 2.2) with the difference that the guard tells the prisoners all cycles, more precisely all members of each cycle (but not the order inside of course).
  Theorem 2.26.
There is no strategy in general.
Proof.
Let $(1,\dots,k_{1})$ be the first cycle created by the guard such that there is $i\in\{1,\dots,k_{1}\}$ such that $p_{i}>0$ and $k_{1}p_{i}>1$. If the guard has created cycles with members less than or equal to $m\in\mbox{$\mathbb{N}$}$, then his next cycle is $(m+1,\dots,m+k_{m})$ such that there is $i\in\{m+1,\dots,m+k_{m}\}$ such that $p_{i}>0$ and $k_{m}p_{i}>1$. Evidently this can be done.
As the prisoners gets altogether 1 as amount, there must be at least one prisoner in each cycle who cannot open all boxes of his cycle. ∎
  Theorem 2.27.
Proof.
There is $m\in\mbox{$\mathbb{N}$}$ such that ${k\sum_{i=m}^{\infty}p_{i}<1}$. If a prisoner’s label is $n\geq m$, and all members of his full cycle $\geq m$, then he gets the amount which is the price of his cycle. Everyone else gets 0. Clearly this strategy works, and the full used amount is less than or equal to ${k\sum_{i=m}^{\infty}p_{i}<1}$. ∎